OpenStudy (anonymous):

What is the length of AB? 14 cm 23.32 cm 26 cm 28.49 cm

OpenStudy (anonymous):

@piglet9

OpenStudy (anonymous):

SOH CAH TOA for the triangle on the left: tan(30) = 10/x xtan(30 = 10 x = 10/tan(30) for the triangle on the right: tan(45) = y/6 y = 6tan(45) where x and y are below: 

OpenStudy (anonymous):

cot 30=AP/10 tan 45=PB/6

OpenStudy (anonymous):

x = 17.33 y = 6 total length = 23.33

OpenStudy (anonymous):

dont forget the cm =D

OpenStudy (anonymous):

thank you both for your help